OpenStudy (anonymous):

12x^(-1/2)+7x^(1/2)+x^3/2. Factor expression completely

OpenStudy (anonymous):

not sure how you are expected to approach this, but if you write it out with the denominator of \[\sqrt{x}\] you will get \[\frac{(x+3)(x+4)}{\sqrt{x}}\]

OpenStudy (anonymous):

oh like this \[x^{-\frac{1}{2}}(12+7x+x^2)\] and then get the answer

OpenStudy (anonymous):

what should the answer be

OpenStudy (anonymous):

?